# Leveraging Multi-source and Heterogeneous Signals for Fatigue Detection

### v2

@@ Appendix A Theoretical Analysis
[58] Suppose the loss function $\mathcal{L}(f(x),y)$ is R-sub-Gaussian under $x\sim\mathbb{P}$ for all $y\in\mathcal{Y}$, then, we have:
@@
Assume that $x,a$ are independent of each other,$x\sim\mathbb{P},a\sim\mathbb{Q}_{a}$, while both being conditionally dependent on variable of $y$ , Let the composite observation be defined as $x^{+}=[x,s],x^{+}\sim\mathbb{P}_{+}$, with the incorporating the variable $a$ , the $\mathcal{E}_{\mathbb{P}_{+}}$ have a tighten upper bound than $\mathcal{E}_{\mathbb{P}}$, the gap $G$ is:
@@
Proof: For ^$\mathcal{E}_{\hat{\mathbb{P}}}\left(f\right)$ and ^$\mathcal{E}_{\hat{\mathbb{P}}_{+}}\left(f\right)$, we have:
@@
Given a random variable $\hat{a}$ generated from a distribution $\mathbb{Q}_{\hat{a}}$, and $\hat{x}^{+}=[x,\hat{a}]\sim\mathbb{Q}$ , then we have:
@@
where ^$\mathcal{E}_{\hat{\mathbb{Q}}}\left(f\right)$ is the empirical generalization error over distribution $\mathbb{Q}$, $d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})$ is the $\mathcal{H}\Delta\mathcal{H}$ distance between $\mathbb{Q}_{a}$ and $\mathbb{Q}_{\hat{a}}$.
@@
$d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{P}_{+},\mathbb{Q})$ $\leq d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{P},\mathbb{P})+d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})$ (22)
$\leq d_{\mathcal{H}\Delta\mathcal{H}}(\mathbb{Q}_{a},\mathbb{Q}_{\hat{a}})$
@@
[59] Let $S$ and $T$ be the source and target domains over $\mathcal{X}\times\mathcal{Y}$, respectively. Let $\mathcal{H}$ be a hypothesis class, and let $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$ be a loss function that is symmetric, obeys the triangle inequality, and is bounded, $\forall(y,y^{\prime})\in\mathcal{Y}^{2}$, $\ell(y,y^{\prime})\leq M$ for some $M>0$. Then, for $h_{S}^{*}=\arg\min\limits_{h\in\mathcal{H}}R_{S}^{\ell}(h)$ and $h_{T}^{*}=\arg\min\limits_{h\in\mathcal{H}}R_{T}^{\ell}(h)$ denoting the ideal hypotheses for the source and target domains, we have

### v3

@@ Appendix A Theoretical Analysis
[bu_tightening_2019] Suppose the loss function $\mathcal{L}(f(x),y)$ is R-sub-Gaussian under $x\sim\mathbb{P}$ for all $y\in\mathcal{Y}$, then, we have:
@@
[…]
@@
Proof: For $\mathcal{E}_{\hat{\mathbb{P}}}\left(f\right)$ and $\mathcal{E}_{\hat{\mathbb{P}}_{+}}\left(f\right)$, we have:
@@
Given a random variable $\hat{a}$ generated from a distribution $\mathbb{Q}_{\hat{a}}$, and $\hat{x}^{+}=[x,\hat{a}$$\sim\mathbb{Q}$ , then we have:
@@
[…]
@@
[…]
@@
[mansour2009domain] Let $S$ and $T$ be the source and target domains over $\mathcal{X}\times\mathcal{Y}$, respectively. Let $\mathcal{H}$ be a hypothesis class, and let $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$ be a loss function that is symmetric, obeys the triangle inequality, and is bounded, $\forall(y,y^{\prime})\in\mathcal{Y}^{2}$, $\ell(y,y^{\prime})\leq M$ for some $M>0$. Then, for $h_{S}^{*}=\arg\min\limits_{h\in\mathcal{H}}R_{S}^{\ell}(h)$ and $h_{T}^{*}=\arg\min\limits_{h\in\mathcal{H}}R_{T}^{\ell}(h)$ denoting the ideal hypotheses for the source and target domains, we have